OpenStudy (anonymous):

Over a two hour time period a snail moved 64 inches. How far is this in yards? Write your answer as a whole number or a mixed number in simplest form. Include the correct unit in your answer.

pooja195 (pooja195):

1 yard = 36 inches

pooja195 (pooja195):

64 inches=? yards We know that 1 yard=36 inches so 64-36=? the remainder you need to make into a decimal then just convert

OpenStudy (anonymous):

36

OpenStudy (anonymous):

no

OpenStudy (anonymous):

28

pooja195 (pooja195):

Wait you can just divide 36/64

OpenStudy (astrophysics):

\[\left( 64 inches \right) \times \left( \frac{ 1 yard }{ 36 inches } \right)\]

OpenStudy (astrophysics):

Note the units cancel out that's all these problems want you to learn, how to cancel units

OpenStudy (astrophysics):

I've cancelled a lot of units in my life...those poor units

OpenStudy (anonymous):

1.7 @pooja195

OpenStudy (calculusxy):

\[64 \div 36 \] Take the remainder and then put it over 36. \(\large \frac{?}{36}\)

pooja195 (pooja195):

You still get 1.7 but as a repeating decimal

OpenStudy (astrophysics):

They want it as a mixed number

OpenStudy (astrophysics):

1+7/9

OpenStudy (astrophysics):

\[1 \frac{ 7 }{ 9 }\]

OpenStudy (calculusxy):

Yes that's correct ^

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Meaning everyone

OpenStudy (astrophysics):

I just hope you know how to get it that way, you can do long division to find the remainder as others have told you

OpenStudy (astrophysics):

I don't think you'll ever use mixed fractions after elementary

pooja195 (pooja195):

I feel like @Astrophysics way of cancelling would have been alot more easier if you aren't good at dividing

OpenStudy (anonymous):

I signed up for this yesterday

pooja195 (pooja195):

Welcome to openstudy :)

OpenStudy (astrophysics):

Oh yes, that method the way I showed with units is how you do it in the sciences and is super convenient all around